the average number of pages in a novel is 276 witha standard deviation of 18 pages. is a sample of 40 novels is randomly chosen, what is the probability that the average number of pages in these books is between 268 and 281?

1
Expert's answer
2022-06-08T18:00:08-0400
"P(268<X<281)=P(Z<\dfrac{281-276}{18/\sqrt{40}})"

"-P(Z\le\dfrac{268-276}{18/\sqrt{40}})"

"=P(Z<1.7568)-P(Z\le-2.8109)"

"=0.96052-0.00247=0.9581"